# SceneForge: Enhancing 3D-text alignment with Structured Scene Compositions

### v1

@@ Abstract
The whole is greater than the sum of its parts—even in 3D-text contrastive learning. We introduce SceneForge, a novel framework that enhances contrastive alignment between 3D point clouds and text through structured multi-object scene compositions. SceneForge leverages individual 3D shapes to construct multi-object scenes with explicit spatial relations, pairing them with coherent multi-object descriptions refined by a large language model. By augmenting contrastive training with these structured, compositional samples, SceneForge effectively addresses the scarcity of large-scale 3D-text datasets, significantly enriching data complexity and diversity. We systematically investigate critical design elements, such as the optimal number of objects per scene, the proportion of compositional samples in training batches, and scene construction strategies. Extensive experiments demonstrate that SceneForge delivers substantial performance gains across multiple tasks, including zero-shot classification on ModelNet, ScanObjNN, Objaverse-LVIS, and ScanNet, as well as few-shot part segmentation on ShapeNetPart. SceneForge’s compositional augmentations are model-agnostic, consistently improving performance across multiple encoder architectures. Moreover, SceneForge improves 3D visual question answering on ScanQA, generalizes robustly to retrieval scenarios with increasing scene complexity, and showcases spatial reasoning capabilities by adapting spatial configurations to align precisely with textual instructions.
@@ 3 Method
We introduce SceneForge (SF), a framework for composing multi‐object 3D scenes by combining individual point clouds with their text descriptions according to explicit spatial relations. SceneForge can be incorporated into any contrastive multimodal 3D–text learning pipeline, and we refer to the resulting approaches as SF-variants. An overview of our pipeline is shown in fig.˜1.
@@ 3.1 SceneForge module
This module constructs a textual description of the combined scene by sequentially incorporating individual object captions $\{t_{0},t_{1},\dots,t_{K-1}\}$ and their corresponding spatial relations $\{s_{1},s_{2},\dots,s_{K-1}\}$. The captioning process mirrors the spatial composition of the 3D scene, starting with the first object’s caption and appending each subsequent caption preceded by its respective spatial relation. However, due to the method’s simplicity, the generated caption may exhibit artifacts such as misplaced punctuation before spatial relations, incorrect capitalization following conjunctions, and disfluent sentence structures. Consequently, we refer to this as a raw caption, $c^{txt,raw}$. To enhance readability and coherence, we refine the raw caption using a large language model (Qwen2.5 [18], obtaining the refined caption $c^{txt}$. The model corrects grammar, punctuation, and structure while preserving the original meaning and spatial relationships. It restructures the text into a fluent, human-like description, splitting overly long sentences when necessary. Beyond improving fluency, this rewriting process also enhances caption diversity and refines OpenShape [13] captions, originally generated by BLIP [11] and Microsoft Azure Cognitive Services (2023), by leveraging the more advanced linguistic capabilities of a recent language model. The full prompt used for refinement is provided in the supplementary, together with an ablaton on the LLM. An overview of the proposed method is provided in fig.˜2.
@@ 3.2 Training Scheme
We instantiate our scene–composition pipeline on three approaches using different point-cloud encoders—OpenShape-PointBERT [13], Uni3D-G [25]- and ViT-Lens-G (a frozen ViT-bigG/14 with trainable adapter lenses) [10] [^1] Uni3D: MIT Licence, OpenShape, ViT-Lens: Apache 2.0 . This variety allows us to verify the encoder-agnostic nature of our method and to select the approach which benefit the most from our pipeline. All variants are trained with their public available code and our modified loss for $200$ epochs with a global batch size of $1152$, $\alpha=0.5$, and a target point-cloud resolution of $P=10\,\text{k}$ points. For caption generation, we use the Qwen 2.5 7B-instruct [18] large language model. SceneForge only requires an additional GPU hosting the lightweight LLM for composition. We pre-generate the first $M$ batches and then switch to a producer–consumer setup: while batch $t$ trains, batch t+Mt\!+\!M is assembled in parallel. Though composition latency is not always fully hidden, parallelization significantly reduce it. For faster models, multiple SceneForge instances and quantization can be used to further amortize overhead. Measured slowdown ranges from 0% to 50% depending on backbone, LLM quantization and replication (see supplementary). For the function $\mathcal{A}^{3D}$, we normalize point clouds to the unit sphere and apply random point dropout and scaling. Rotation and translation strategies differ based on whether we process single objects, objects intended for combination in a scene, or the final composed scene. For single objects, we allow both shifts and larger rotations. However, for objects that will be combined, shifts are disabled to avoid inconsistencies in composition, while full rotations around the vertical axis and slight rotations along other axes are permitted. The latter ensures that concepts such as “over” and “under” remain semantically meaningful. For the final combined point cloud, we adopt the same rotation constraints as in the previous case to preserve spatial semantics but additionally allow translations.
@@ 4.1 Zero-Shot Classification
We follow the standard zero-shot evaluation protocol on Objaverse LVIS, ModelNet40 and ScanObjNN, where categories are mapped to text prompts by formatting a set of templates (e.g., “a point cloud model of a ,” “There is a in the scene.”) and the model is evaluated on the classification accuracy. Additonally, adopting the pipeline from $\operatorname{CLIP}^{2}$ [21], we test our models on the Scannet [4] dataset to evaluate their zero-shot performance on object instances from real-world scenarios.
@@
Figure˜3 tracks zero-shot accuracy for three contrastive learners—OpenShape-PointBERT [13], Uni3D (EVA-Giant) [25, 5], and ViT-Lens-G [10]—as we vary the maximum number of shapes that SceneForge can merge during training. A consistent trend emerges. From N=1N{=}1 to N=3N{=}3, accuracy rises monotonically: LVIS gains +0.81.3+0.8-1.3 pp, ScanObjNN +0.81.3+0.8-1.3 pp, ScanNet +1.43.6+1.4-3.6 pp, and ModelNet40 about +0.8+0.8 pp. Improvements peak at N=3N{=}3; Uni3D benefits most, while the lighter OpenShape also advances. ViT-Lens shows a more modest gain, plausibly because its frozen CLIP backbone trains only lightweight adapters, offering less plasticity when confronted with composed shapes. Increasing to N=4N{=}4 plateaus on canonical datasets and already trims accuracy on LVIS and ScanNet, whose higher intra-class variability makes them clutter-sensitive. At N=5N{=}5 the drop is universal, indicating that squeezing five objects into a fixed 10 k-point budget fragments salient geometry and introduces caption noise, hampering alignment. Interestingly, performance variations are relative minor on ModelNet40 and ScanObjNN, but the effect is far more pronounced on LVIS and ScanNet. We attribute this to the latter’s more complex object distributions: as additional shape combinations are introduced, the model must balance greater intra-class variability with the need for discriminative features, a trade-off that becomes harder to resolve in these richer, noisier datasets. Considering these results, we adopt N=3N{=}3 as default SF-variants and report the full sweeps (N=15)(N{=}1\!-\!5) in the supplementary.
@@
Effect of SceneForge. Across all three backbones SceneForge delivers consistent top-1 gains. With the “no-LVIS” training split the average improvement is +1.50+1.50 pp for OpenShape, +0.78+0.78 pp for ViT-Lens and +1.73+1.73 pp for Uni3D; with the “+LVIS” split the corresponding gains are +1.43+1.43, +0.85+0.85 and +1.75+1.75 pp. Uni3D benefits most—likely a consequence of its larger capacity, which can better exploit the richer intra-sample diversity injected by multi-shape compositions—yet even the smaller OpenShape and the adapter-based ViT-Lens improve.
@@
Why do multi-object compositions enhance single-object classification? While it might initially appear counterintuitive, training with multi-object compositions fosters improved representations that benefit single-object recognition tasks. This phenomenon aligns with established findings in representation learning literature, particularly in image classification, where multi-sample augmentations (e.g., CutMix, MixUp) are known to induce smoother decision boundaries and promote robust generalization by exposing models to more diverse feature combinations. Analogously, in our structured 3D scene compositions, the increased complexity and relational context implicitly regularize the learned representations, facilitating the emergence of discriminative features resilient to variations in single-object scenarios encountered at inference.
@@ 4.2 Few-Shot Part Segmentation on PartNet
We assess how our multi‐shape pretraining enhances 3D–text integration in the ScanQA benchmark [7], which requires answering natural‐language questions about ScanNet scenes. Following prior works, we freeze each 3D encoder and attach it to BLIP2–FlanT5 [3, 12], then fine‐tune on ScanQA’s question–answer pairs. In addition to the sota contrastive baselines (OmniBind-Large, OmniBind-Full), we include SF‐OpenShape, SF‐ViT-Lens and SF-Uni3D as our multi‐shape variants, alongside their single‐shape counterparts. We report here BLEU-4 (B-4), CIDEr and Exact Match (EM); full metrics appear in the supplementary. Table˜3 gives the results.
All three backbones see substantial gains with SceneForge: SF-OpenShape improves B-4 by +1.8 pp, CIDEr by +6.7 pp and EM by +2.8 pp; SF-ViT-Lens adds +1.3 pp, +5.9 pp and +2.1 pp, respectively; and SF-Uni3D leads with +2.9 pp, +8.4 pp and +4.1 pp. While all backbones gain, Uni3D yields the largest relative and absolute boosts—consistent with its greater capacity—whereas ViT-Lens, despite smaller gains, still surpasses OmniBind. A qualitative review shows that, while baseline encoders match our variants on attribute- or color-based questions, SceneForge variants significantly outperform them on spatial reasoning queries—e.g. “What is over the brown chair?”—where modeling inter-object relationships is essential. This suggests that our structured multi-shape augmentation not only sharpens local feature representations but also boosts the encoder’s ability to infer complex spatial configurations, a critical capability for 3D scene understanding.
@@ 4.4 N-Objects cross-modal Retrieval
Following previous works [17, 25], we evaluate our approach on cross-modal retrieval. Specifically, we consider models trained on the ensemble dataset excluding Objaverse-LVIS and evaluate both 3D-to-text and text-to-3D retrieval on this unseen dataset. To assess retrieval performance, we embed all shapes using the model and compute the cosine similarity between samples across the two modalities. The retrieval process ranks samples based on similarity, and we compute top-k accuracy metrics. Beyond standard single-object retrieval, we aim to analyze how well our models understand multi-object scenes. To this end, we introduce the "N-LVIS" dataset, where each object is combined with N1N-1 additional objects, increasing scene complexity.
For ease of visualization, we report here only results for the best SceneForge backbone, Uni3D ($N=1$) and all its SF-Uni3D variants ($N\in\{2..5\}$), and we compare to the state-of-the-art OmniBind-F [16] model. The same analysis for the other SceneForge backbones is deferred to the supplementary, as they exhibit similar gains and trends.
We evaluate the retrieval performance on N-LVIS from $N=1$ (standard cross-modal retrieval) to $N=10$. We plot in in fig.˜4 the averaged top-1 accuracy (text to 3D and viceversa) for the considered methods. As expected, prior models struggle when faced with multi-object compositions, with top-1 accuracy falling below $50\%$ at just $N=2$. In contrast, our models, trained on varying levels of multi-object compositions, show a clear generalization advantage. Notably, each of our models peaks at its respective maximum number of training shapes. More importantly, we show that training on $N>1$ significantly improves retrieval at higher N, demonstrating that learning from compositions enables better understanding of complex multi-object scenes. Our N=3 model, for instance, maintains an over $70\%$ accuracy even at $N=6$ and roughly $60\%$ at $N=7$. This generalization comes with a trade-off: as $N$ increases, models trained on more complex compositions show slightly weaker performance at lower $N$. This remarks the findings in section˜4.1 that training on simpler compositions better preserves single-object understanding, while showing that multi-object training encourages broader compositional understanding. Extended Top-1 and Top-5 results are reported in the supplementary.
@@ 4.5 Object Repositioning
To qualitatively assess our approach ability to improve reasoning about spatial relationships, we introduce a simple object repositioning task. Given two objects, we combine them in an initial configuration using our MSF. The combined caption is then modified to describe a new spatial relationship, and we optimize a three-parameter offset vector to reposition the second object, maximizing alignment with the updated description while keeping the 3D-text encoder frozen. Full optimisation details appear in the supplementary. As illustrated in fig.˜5 encoders trained with SceneForge reliably relocate the object to the target position, while baseline encoders stall in semantically plausible yet misaligned configurations. We observe the same qualitative behaviour with all other variants, confirming that SceneForge strengthens spatial and compositional reasoning.

### v2

@@ Abstract
The whole is greater than the sum of its parts, even in 3D-text contrastive learning. We introduce SceneForge, a novel framework that enhances contrastive alignment between 3D point clouds and text through structured multi-object scene compositions. SceneForge leverages individual 3D shapes to construct multi-object scenes with explicit spatial relations, pairing them with coherent multi-object descriptions refined by a large language model. By augmenting contrastive training with these structured, compositional samples, SceneForge effectively addresses the scarcity of large-scale 3D-text datasets, significantly enriching data complexity and diversity. We systematically investigate critical design elements, such as the optimal number of objects per scene, the proportion of compositional samples in training batches, and scene construction strategies. Extensive experiments demonstrate that SceneForge delivers substantial performance gains across multiple tasks, including zero-shot classification on ModelNet, ScanObjNN, Objaverse-LVIS, and ScanNet, as well as few-shot part segmentation on ShapeNetPart. SceneForge’s compositional augmentations are model-agnostic, consistently improving performance across multiple encoder architectures. Moreover, SceneForge improves 3D visual question answering on ScanQA, generalizes robustly to retrieval scenarios with increasing scene complexity, and showcases spatial reasoning capabilities by adapting spatial configurations to align precisely with textual instructions.
@@ 3 Method
We introduce SceneForge (SF), a framework for composing multi‐object 3D scenes by combining individual point clouds with their text descriptions according to explicit spatial relations. SceneForge can be incorporated into any contrastive multimodal 3D–text learning pipeline, and we refer to the resulting approaches as SF-variants. An overview of our pipeline is shown in Figure˜1.
@@ 3.1 SceneForge module
This module constructs a textual description of the combined scene by sequentially incorporating individual object captions $\{t_{0},t_{1},\dots,t_{K-1}\}$ and their corresponding spatial relations $\{s_{1},s_{2},\dots,s_{K-1}\}$. The captioning process mirrors the spatial composition of the 3D scene, starting with the first object’s caption and appending each subsequent caption preceded by its respective spatial relation. However, due to the method’s simplicity, the generated caption may exhibit artifacts such as misplaced punctuation before spatial relations, incorrect capitalization following conjunctions, and disfluent sentence structures. Consequently, we refer to this as a raw caption, $c^{txt,raw}$. To enhance readability and coherence, we refine the raw caption using a large language model (Qwen2.5 [24]), obtaining the refined caption $c^{txt}$. The model corrects grammar, punctuation, and structure while preserving the original meaning and spatial relationships. It restructures the text into a fluent, human-like description, splitting overly long sentences when necessary. Beyond improving fluency, this rewriting process also enhances caption diversity and refines OpenShape [17] captions, originally generated by BLIP [14] and Microsoft Azure Cognitive Services (2023), by leveraging the more advanced linguistic capabilities of a recent language model. The full prompt used for refinement is provided in the supplementary, together with an ablaton on the LLM. An overview of the proposed method is provided in Figure˜2.
@@ 3.2 Training Scheme
We instantiate our scene–composition pipeline on three approaches using different point-cloud encoders: OpenShape-PointBERT [17], Uni3D-G [31], and ViT-Lens-G (a frozen ViT-bigG/14 with trainable adapter lenses) [13]. This variety allows us to verify the encoder-agnostic nature of our method and to select the approach which benefit the most from our pipeline. All variants are trained with their public available code and our modified loss for $200$ epochs with a global batch size of $1152$, $\alpha=0.5$, and a target point-cloud resolution of $P=10\,\text{k}$ points. This point budget was chosen as a trade-off between detail and efficiency, as detailed in the supplementary. For caption generation, we use the Qwen 2.5 7B-instruct [24] large language model. SceneForge only requires an additional GPU hosting the lightweight LLM for composition. We pre-generate the first $M$ batches and then switch to a producer–consumer setup: while batch $t$ trains, batch t+Mt\!+\!M is assembled in parallel. Though composition latency is not always fully hidden, parallelization significantly reduce it. For faster models, multiple SceneForge instances and quantization can be used to further amortize overhead. Measured slowdown ranges from 0% to 50% depending on backbone, LLM quantization and replication (see supplementary). For the function $\mathcal{A}^{3D}$, we normalize point clouds to the unit sphere and apply random point dropout and scaling. Rotation and translation strategies differ based on whether we process single objects, objects intended for combination in a scene, or the final composed scene. For single objects, we allow both shifts and larger rotations. However, for objects that will be combined, shifts are disabled to avoid inconsistencies in composition, while full rotations around the vertical axis and slight rotations along other axes are permitted. The latter ensures that concepts such as “over” and “under” remain semantically meaningful. For the final combined point cloud, we adopt the same rotation constraints as in the previous case to preserve spatial semantics but additionally allow translations.
@@ 4.1 Zero-Shot Classification
We follow the standard zero-shot evaluation protocol on Objaverse LVIS [6], ModelNet40 [22] and ScanObjNN [20], where categories are mapped to text prompts by formatting a set of templates (e.g., “a point cloud model of a ) and the model is evaluated on the classification accuracy. Additonally, adopting the pipeline from $\operatorname{CLIP}^{2}$ [27], we test our models on the Scannet [5] dataset to evaluate their zero-shot performance on object instances from real-world scenarios.
@@
Figure˜3 tracks zero-shot accuracy for three contrastive learners: OpenShape-PointBERT [17], Uni3D (EVA-Giant) [31, 7], and ViT-Lens-G [13], as we vary the maximum number of shapes that SceneForge can merge during training. A consistent trend emerges. From N=1N{=}1 to N=3N{=}3, accuracy rises monotonically: LVIS gains +0.81.3+0.8-1.3 pp, ScanObjNN +0.81.3+0.8-1.3 pp, ScanNet +1.43.6+1.4-3.6 pp, and ModelNet40 about +0.8+0.8 pp. Improvements peak at N=3N{=}3; Uni3D benefits most, while the lighter OpenShape also advances. ViT-Lens shows a more modest gain, plausibly because its frozen CLIP backbone trains only lightweight adapters, offering less plasticity when confronted with composed shapes. Increasing to N=4N{=}4 plateaus on canonical datasets and already trims accuracy on LVIS and ScanNet, whose higher intra-class variability makes them clutter-sensitive. At N=5N{=}5 the drop is universal, indicating that squeezing five objects into a fixed 10 k-point budget fragments salient geometry and introduces caption noise, hampering alignment. Interestingly, performance variations are relative minor on ModelNet40 and ScanObjNN, but the effect is far more pronounced on LVIS and ScanNet. We attribute this to the latter’s more complex object distributions: as additional shape combinations are introduced, the model must balance greater intra-class variability with the need for discriminative features, a trade-off that becomes harder to resolve in these richer, noisier datasets. Considering these results, we adopt N=3N{=}3 as default SF-variants and report the full sweeps (N=15)(N{=}1\!-\!5) in the supplementary.
@@
Effect of SceneForge. Across all three backbones SceneForge delivers consistent top-1 gains. With the “no-LVIS” training split the average improvement is +1.50+1.50 pp for OpenShape, +0.78+0.78 pp for ViT-Lens and +1.73+1.73 pp for Uni3D; with the “+LVIS” split the corresponding gains are +1.43+1.43, +0.85+0.85 and +1.75+1.75 pp. Uni3D benefits most, likely a consequence of its larger capacity, which can better exploit the richer intra-sample diversity injected by multi-shape compositions, yet even the smaller OpenShape and the adapter-based ViT-Lens improve.
@@
Why do multi-object compositions enhance single-object classification? While it might initially appear counterintuitive, training with multi-object compositions fosters improved representations that benefit single-object recognition tasks. This phenomenon aligns with established findings in representation learning literature, particularly in image classification, where multi-sample augmentations (e.g., CutMix, MixUp) are known to induce smoother decision boundaries and promote robust generalization by exposing models to more diverse feature combinations. Analogously, in our structured 3D scene compositions, the increased complexity and relational context implicitly regularize the learned representations, facilitating the emergence of discriminative features resilient to variations in single-object scenarios encountered at inference. By analyzing the positive–negative similarity margins in zero-shot classification, we find that SF-variants yield larger margins over baselines, indicating stronger inter-class separation and more robust decision boundaries. We report the quantitative analysis in the supplementary material.
@@ 4.2 Few-Shot Part Segmentation on PartNet
We further assess 3D–text alignment on the ScanQA benchmark [2, 9], which requires answering natural‐language questions about ScanNet scenes. Following prior works, we freeze each 3D encoder and attach it to BLIP2–FlanT5 [4, 15], then fine‐tune on ScanQA’s question–answer pairs. In addition to the sota contrastive baselines (OmniBind-Large, OmniBind-Full), we include SF‐OpenShape, SF‐ViT-Lens and SF-Uni3D as our multi‐shape variants, alongside their single‐shape counterparts. We report here BLEU-4 (B-4), CIDEr and Exact Match (EM); full metrics appear in the supplementary. Table˜3 gives the results. All three backbones see substantial gains with SceneForge: SF-OpenShape improves B-4 by +1.8 pp, CIDEr by +6.7 pp and EM by +2.8 pp; SF-ViT-Lens adds +1.3 pp, +5.9 pp and +2.1 pp, respectively; and SF-Uni3D leads with +2.9 pp, +8.4 pp and +4.1 pp. While all backbones gain, Uni3D yields the largest relative and absolute boosts, consistent with its greater capacity, whereas ViT-Lens, despite smaller gains, still surpasses OmniBind. A qualitative review shows that, while baseline encoders match our variants on attribute- or color-based questions, SceneForge variants significantly outperform them on spatial reasoning queries, e.g. “What is over the brown chair?”, where modeling inter-object relationships is essential. This suggests that our structured multi-shape augmentation not only sharpens local feature representations but also boosts the encoder’s ability to infer complex spatial configurations, a critical capability for 3D scene understanding.
4.4 Supervised Fine-Tuning
To test if the benefits of SCENEFORGE pre-training extend to supervised settings, we evaluate full fine-tuning and Parameter-Efficient Fine-Tuning (PEFT) methods [10, 32, 16]. We perform experiments on three of the considered classification benchmarks: the synthetic ModelNet40, the challenging real-world ScanObjectNN, and ScanNet Instances extracted from complex indoor scenes.
Table 4: Supervised fine-tuning accuracy (%).
Model Method Trainable Params ModelNet40 ScanObjectNN ScanNet Inst.
Uni3D Full Fine-Tuning 1016.5M (100%) 94.28 97.12 82.72
Adapter 7.6M (0.74%) 94.35 96.80 81.42
DAPT 7.3M (0.72%) 94.33 96.78 82.65
PointGST 4.1M (0.40%) 94.83 97.68 83.04
SF-Uni3D Full Fine-Tuning 1016.5M (100%) 94.42 97.58 83.58
Adapter 7.6M (0.74%) 94.46 97.09 82.56
DAPT 7.3M (0.72%) 94.49 97.15 83.46
PointGST 4.1M (0.40%) 94.95 98.09 84.29
As shown in Table 4, our SF-Uni3D backbone provides a superior initialization, consistently outperforming the baseline, especially on the complex ScanNet Instances. A strong synergy with PEFT is also evident: the PointGST method surpasses even full fine-tuning while using just 0.4% of the trainable parameters.
@@ 4.5 N-Objects cross-modal Retrieval
Following prior work [23, 31], we evaluate cross-modal retrieval on the unseen Objaverse-LVIS dataset, measuring top-k accuracy for both 3D-to-text and text-to-3D retrieval. Models are trained on the full ensemble excluding this set, and performance is assessed via cosine similarity retrieval across embedded samples. Beyond standard single-object retrieval, we aim to analyze how well our models understand increasingly multi-object scenes. To this end, we introduce the N-LVIS benchmark, where each shape is composed with N1N-1 additional objects. For clarity, we report results only for the best SF backbone (Uni3D with $N=1..5$) and compare against OmniBind-F [21]; analogous results for other backbones are provided in the supplementary.
We evaluate retrieval on N-LVIS from N=1N{=}1 (standard retrieval) to N=10N{=}10, reporting averaged top-1 accuracy (text to 3D and viceversa) in Figure˜4. Prior models degrade sharply when faced with multi-object compositions, dropping below 50% at N=2N{=}2. In contrast, our models, trained with varying numbers of composed objects, exhibit strong generalization, each peaking near its training composition size. Notably, the N=3N{=}3 model sustains over 70% accuracy at N=6N{=}6 and around 60% at N=7N{=}7, highlighting its robustness in complex scenes. This comes at a slight cost: models trained with higher $N$ perform marginally worse at lower $N$, reinforcing our findings in Section˜4.1 that simpler compositions better preserve single-object understanding. Full top-1 and top-5 results are provided in the supplementary.
@@ 4.6 Object Repositioning
To qualitatively assess our approach ability to improve reasoning about spatial relationships, we introduce a simple object repositioning task. Given two objects, we combine them in an initial configuration using our MSF. The combined caption is then modified to describe a new spatial relationship, and we optimize a three-parameter offset vector to reposition the second object, maximizing alignment with the updated description while keeping the 3D-text encoder frozen. Full optimisation details appear in the supplementary. As illustrated in Figure˜5 encoders trained with SceneForge reliably relocate the object to the target position, while baseline encoders stall in semantically plausible yet misaligned configurations. We observe the same qualitative behaviour with all other variants, confirming that SceneForge strengthens spatial and compositional reasoning.